what is the equation for a line containing the point (9,-9) with a slope of 6/7 ?
Answer by nerdybill(7384)

what is the equation for a line containing the point (9,-9) with a slope of 6/7 ?
Plug given data into the "point-slope" form:
y - y1 = m(x - x1)
y - (-9) = (6/7)(x - 9)
y + 9 = (6/7)x - 54/7
y = (6/7)x - 54/7 - 9
y = (6/7)x - 54/7 - 63/7
y = (6/7)x - 117/7 (this is in "slope-intercept" form)